A one litre box of juice is three quarters full.you pour out one fifth of the litre of juice.how much of the litre of juice is now left in the box


1
Expert's answer
2022-01-07T12:35:24-0500

"\frac{3}{4}-\frac{1}{5}=\frac{15}{20}-\frac{4}{20}=\frac{11}{20}"


Answer: 11/20 of the litre of juice is left in the box (it is 550 ml).